Find the slope of the line that passes through the pair of points
(1,7), (10,1)

Answer by Alan3354(69443)

Find the slope of the line that passes through the pair of points
(1,7), (10,1)
-------------
Slope m = (y2-y1)/(x2-x1)
m = (1-7)/(10-1)
m = -6/9
m = -2/3
-----------
Always use points left to right, or increasing value of x.